# CFSP: An Efficient Structured Pruning Framework for LLMs with Coarse-to-Fine Activation Information

### v1

@@ B.1 More Results on LLaMA
In addition to the experiments presented in Section 3.2, we also conduct experiments on the latest powerful model LLaMA3.1-8B. The results are shown in Table 8. In the single-shot pruning setting, CFSP consistently outperforms other baselines across a variety of tasks and sparsity budgets. Furthermore, we experiment with recovery fine-tuning for CFSP and Wanda-SP. As observed with previous models, our method still achieves better results.
@@ B.2 More Analysis of CFSP
In Equation 3, we introduce a hyperparameter $\alpha$ to control the intensity of significance during calculating block importance. In preliminary experiments, we explore the impact of different $\alpha$ and the results are shown in Figure 9. We observe that for smaller models like LLaMA3-8B, a smaller $\alpha$ is better, while for larger model like LLaMA3-70B model, a larger $\mathbf{\alpha}$ is more appropriate. Finally, we set $\alpha=1$ for the 7B/8B/13B models and $\alpha=3$ for the 70B models.
@@
As described in Appendix A.2, we explore various datasets for recovery fine-tuning. As shown in Table 12, FineWeb-Edu consistently outperforms others in a variety of tasks, particularly demonstrating significant improvements in knowledge-intensive tasks such as MMLU and FreebaseQA, which is shown challenging for pruned models (Jaiswal et al., 2023a). Thus, we select it for recovery fine-tuning.

### v2

@@ B.1 More Results on LLaMA
In addition to the experiments presented in Section 3.2, we also conduct experiments on the latest powerful model LLaMA3.1-8B.
The results are shown in Table 9. In the single-shot pruning setting, CFSP consistently outperforms other baselines across a variety of tasks and sparsity budgets. Furthermore, we experiment with recovery fine-tuning for CFSP and Wanda-SP. As observed with previous models, our method still achieves better results.
@@
Results of Qwen2.5
In addition to the models of the LLaMA families, we also conduct experiments on Qwen2.5-7B (Qwen Team, 2024) to verify whether our pruning framework is model-agnostic. As shown in Table 8, CFSP consistently shows better zero-shot performance on average and achieves comparable PPL.
@@ B.2 More Analysis of CFSP
In Equation 3, we introduce a hyperparameter $\alpha$ to control the intensity of significance during calculating block importance. In preliminary experiments, we explore the impact of different $\alpha$ and the results are shown in Figure 9. We observe that for smaller models like LLaMA3-8B, a smaller $\alpha$ is better, while for larger models like the LLaMA3-70B model, a larger $\mathbf{\alpha}$ is more appropriate. Finally, we set $\alpha=1$ for the 7B/8B/13B models and $\alpha=3$ for the 70B models.
@@
As described in Appendix A.2, we explore various datasets for recovery fine-tuning. As shown in Table 13, FineWeb-Edu consistently outperforms others in a variety of tasks, particularly demonstrating significant improvements in knowledge-intensive tasks such as MMLU and FreebaseQA, which is shown challenging for pruned models (Jaiswal et al., 2024). Thus, we select it for recovery fine-tuning.